OpenStudy (anonymous):

What is the sum of the arithmetic sequence 135,123,111,... if there are 34 terms? Answers to choose from. a)-2,176 b)-2,142 c)-2,108 d)-2,074

OpenStudy (anonymous):

what is the common difference?

OpenStudy (anonymous):

idk

OpenStudy (anonymous):

the common difference of an AP is the difference between each term. d = (n-1)term-(n)term

OpenStudy (anonymous):

the difference between two consecutive terms is the common difference. What is it?

OpenStudy (anonymous):

imreally not sure i dont understand any of this can you guide me though it

OpenStudy (anonymous):

lets say i have an AP of 5 terms 1,5,9,13,17 the common difference: d = 5-1 = 9-5 = 13-9 = 17-13 = 4

OpenStudy (anonymous):

so what is the common difference for your AP?

OpenStudy (anonymous):

7

OpenStudy (anonymous):

how did you get 7?

OpenStudy (anonymous):

I subtracted all the numbers i had like you did

OpenStudy (anonymous):

i did not subtract all the numbers.... 5-1 = 4 9 - 5 = 4 13 - 9 = 4 17 - 13 = 4 as you can see, the common difference is constant since it is an AP and it is 4. I did not subtract all the numbers. I only subtracted two consecutive numbers

OpenStudy (anonymous):

it does not matter which terms you choose as long as those two terms are consecutive

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

i got 7 again

OpenStudy (anonymous):

I got a final answer of -2,176 is this correct??

OpenStudy (anonymous):

no it is not. I seriously wonder if you are even trying at all... I practically gave you the way to find the common difference of the question and you still can give the wrong answer...

OpenStudy (anonymous):

d = 123 - 135 = -12 or d = 111 - 123 = -12 how in the world do you keep getting 7?

OpenStudy (anonymous):

i plugged in every thing you have said and my calculator keeps giving me the answers im giving you :(

OpenStudy (anonymous):

@hartnn

OpenStudy (anonymous):

could you please tell me if i have the correct answer? -2,176

hartnn (hartnn):

i think that incorrect, what did u get d as ?

OpenStudy (anonymous):

originally 7 but irkiz helped me and gave me -12

hartnn (hartnn):

yes, d=-12 S=34/2 (2*135+33*(-12)) = ?

OpenStudy (anonymous):

Okay thank you

OpenStudy (anonymous):

oh -2,142 thankyou

hartnn (hartnn):

yes. welcome ^_^